use the pythagorean theorem to find the indicated quantity. Round your answer to the nearest hundredth.
14^2 = 5^2 + a^2

Answer by stanbon(75887)

Round your answer to the nearest hundredth.
14^2 = 5^2 + a^2
a^2 = 196 - 25
a^2 = 171
a = sqrt(171)
a = 13.077
===================
Cheers,
Stan H.